### v4

# Hybrid Swin Attention Networks for Simultaneously Low-Dose PET and CT Denoising

Abstract
Low-dose computed tomography (LDCT) and positron emission tomography (PET) have emerged as safer alternatives to conventional imaging modalities by significantly reducing radiation exposure. However, this reduction often results in increased noise and artifacts, which can compromise diagnostic accuracy. Consequently, denoising for LDCT/PET has become a vital area of research aimed at enhancing image quality while maintaining radiation safety. In this study, we introduce a novel Hybrid Swin Attention Network (HSANet), which incorporates Efficient Global Attention (EGA) modules and a hybrid upsampling module. The EGA modules enhance both spatial and channel-wise interaction, improving the network’s capacity to capture relevant features, while the hybrid upsampling module mitigates the risk of overfitting to noise. We validate the proposed approach using a publicly available LDCT/PET dataset. Experimental results demonstrate that HSANet achieves superior denoising performance compared to existing methods, while maintaining a lightweight model size suitable for deployment on GPUs with standard memory configurations. This makes our approach highly practical for real-world clinical applications.
keywords:
Image denoising , Swin Transformer , low-dose CT , low-dose PET
journal: Computerized Medical Imaging and Graphics
[…]
[label1]organization=IWR, Heidelberg University,city=Heidelberg, postcode=69120, state=Baden Württemberg, country=Germany
[…]
1 Introduction
Medical imaging modalities such as computed tomography (CT) and positron emission tomography (PET) are indispensable tools for the diagnosis, staging, and monitoring of a wide range of diseases. However, repeated or high-dose exposure to ionizing radiation poses significant health risks to patients, particularly in routine screening, pediatric imaging, or longitudinal follow-up [1, 2]. To mitigate these risks, low-dose imaging protocols have been widely adopted [3, 4]. While effective in reducing radiation exposure, these protocols often lead to images with increased noise, reduced contrast, and compromised diagnostic quality [5, 6, 7]. Consequently, developing robust denoising techniques that can preserve anatomical and functional fidelity while suppressing noise has become a critical area of research. Recent advances in deep learning and data-driven reconstruction methods offer promising solutions to this challenge, enabling high-quality image recovery from low-dose acquisitions across CT and PET modalities [8, 9, 10, 11].
Especially, with deep learning has dominated image denoising research due to their spatial locality and efficiency [12, 13, 14], lots of works have been applied in low-dose CT (LDCT) and LDPET denoising [15, 10, 16]. Chen et al. [10] combines convolutional autoencoder architecture with residual learning to effectively suppress noise in LDCT images while preserving structural details. Li et al. [16] explore unpaired deep learning approaches for LDCT denoising, demonstrating that generative adversarial network (GAN) such as CycleGAN can effectively learn noise reduction mappings from unpaired normal-dose and LDCT images without requiring matched training data. However, their ability to capture long-range dependencies is inherently limited [17]. This shortcoming motivates the exploration of alternative architectures capable of modeling global relationships. Transformer-based architectures, originally developed for natural language processing [18], have recently shown remarkable success in various computer vision tasks by leveraging self-attention mechanisms to model global context [19]. Wang et al. [11] introduce a fully convolution-free architecture for LDCT denoising, leveraging a Token2Token dilated vision transformer to effectively model both local textures and global anatomical structures through hierarchical self-attention. While Vision Transformers (ViTs) demonstrate strong global modeling capabilities through full self-attention, their fixed patch structure and high computational cost limit their effectiveness in dense prediction tasks such as CT denoising. To overcome these issues, Swin Transformer [20] introduces a hierarchical, shifted-window self-attention mechanism that enables both scalability and improved local-context modeling. Jian et al. [21] leverages shifted window attention to effectively balance global context modeling with local structural preservation for LDCT denoising. Zhu et al. [22] incorporates a noise extraction subnetwork and hierarchical multi-scale attention to effectively suppress noise and artifacts while preserving structural detail and contrast. In addition, the adaptation of Swin Transformer for image restoration was further refined in SwinIR [23], which integrates residual connections and a shallow feature extraction module into the Swin framework, making it well-suited for denoising applications. Wang et al. [24] propose a self-pretraining strategy for LDCT denoising by adapting masked autoencoders within a SwinIR-based architecture, enabling effective structural preservation and noise suppression. Wang et al. [25] present a comprehensive evaluation of five state of the art AI models for restoring low-count whole-body PET//MRI images across a wide dose reduction spectrum, demonstrating that SwinIR consistently achieves superior quantitative and qualitative performance, particularly at clinically significant ultra-low-dose levels. While SwinIR leverages supervised learning with hierarchical attention to generalize across datasets, Deep Image Prior (DIP) [26, 27] offers a complementary architecture level design that upsampling can significantly affect denoising performance. Gong et al. [28] pioneered the application of Deep Image Prior (DIP) to static PET image reconstruction, proposing a personalized, unsupervised framework that leverages patient-specific prior information without requiring large training datasets. Further on, Sun et al. [29] combine DIP with regularization by denoising to enhance quantitative accuracy that requires no ground truth data and optionally accepts prior static PET images, outperforming traditional denoising methods across both simulated and clinical datasets.
In this study, we proposed a deep learning model for LDCT/PET image denoising, called the Hybrid Swin Attention Network (HSANet). The architecture is inspired by SwinIR and Swin-Unet, leveraging their ability to capture both local and long-range dependencies, as well as the hierarchical structure of medical images. To further enhance generalization and robustness, we introduced a novel patch expanding block, which focuses on learning from low-frequency textures, helping the model avoid overfitting to high-frequency noise. In addition, we improved the global attention mechanism by drastically reducing its parameter count, making it significantly more efficient. Our experiments demonstrate that HSANet outperforms mainstream denoising methods, achieving superior denoising performance with fewer parameters and lower GPU memory consumption, making it highly suitable for practical applications.
The main contributions of the papers are as follows:
  • 1.
    We propose a novel LDCT/PET denoising model, HSANet, which combines the strengths of both convolutional networks and transformers to effectively capture local and long-range dependencies. This hybrid design leads to more stable optimization and consistently improved denoising results.
  • 2.
    We propose a Hybrid Interpolation Convolution (HIC) patch expanding block, which encourages the model to focus on low-frequency image structures, helping to suppress overfitting to high-frequency noise.
  • 3.
    We propose an attention module, EGA, which comprises two variants: the Efficient Sequential Global Attention (ESGA) and Efficient Parallel Global Attention (EPGA). ESGA is applied at the skip connections, while EPGA is embedded within the Swin Transformer blocks. Together, they enhance information fusion between the encoder and decoder, suppress noise, and achieve higher Peak Signal-to-Noise Ratio (PSNR) with significantly fewer parameters.
2 Methods
In this section, we will introduce HSANet, ESGA and EPGA module, as well as HIC patch expand block, in the following subsections.
2.1 HSANet
The architecture of HSANet (illustrated in Fig. LABEL:HSANet) consists of two main components: a convolutional branch and a Swin Transformer-based branch. The Swin Transformer blocks are responsible for capturing long-range dependencies within LDCT images. Inspired by Swin-Unet [20], we adopt a hierarchical encoder-decoder structure to effectively extract multi-scale features. To capture local dependencies, we have adopted residual convolutional blocks from ResNet, which proved more effective and stable. Additionally, we have incorporated a convolutional layer with residual connections outside the encoder-decoder path, following the design of SwinIR. To enhance texture preservation and edge sharpness, we have used a residual multi-layer perceptron (MLP) structure, which helps retain high-frequency details while enabling effective feature fusion. We have proposed two types of Efficient Global Attention modules, ESGA and EPGA. ESGA is applied within the residual convolution blocks, while EPGA is used in the decoder. We have also replaced MLP in Swin Transformer block with ESGA. The proposed attention modules are used for richer feature interaction and to suppress redundant or noisy information. Details of these modules are provided in Section 2.2. Finally, we have proposed a patch expanding module in the decoder, designed to learn low-frequency structures such as background tissue, organs, and bones, instead of focusing on high-frequency noise. Further details are provided in Section 2.3.
2.2 Attention module
2.2.1 ESGA module
Liu et al. [30] demonstrated significant performance improvements over the CBAM [31] and BAM [32] modules. However, the number of parameters increases significantly as the model size grows. To address this, we replace the original $7\times 7$ convolution kernel with a more lightweight $1\times 1$ convolution, effectively reducing the number of parameters. To compensate for the reduced receptive field and to capture broader spatial patterns and contextual features, we adopt an (inverse) pixel shuffle operation as proposed in [33]. Fig. LABEL:ESGAa shows the overview structure. Specifically, let the input feature be $F_{1}\in\mathbb{R}^{N\times H\times W\times C}$. For channel attention, we employ an encoder-decoder MLP structure to enhance channel-wise information interaction (Fig. LABEL:ESGAb), which is formulated as follows:
$F_{channel}=F_{1}\circledast sigmoid(L_{2}(L_{1}(F_{1})))$ (1)
[…]
For spatial attention, an encoder-decoder $1\times 1$ convolution structure is used to improve spatial-wise information interaction (Fig. LABEL:ESGAc), which can be represented as below:
$F_{2}=F_{channel}\circledast sigmoid(\pi^{-1}(\mathcal{PS}(Conv_{2}(Conv_{1}(\pi(\mathcal{PS}^{-1}(F_{channel})))))))$ (2)
where $\mathcal{PS}$ is periodic shuffling operator. More detailed information can be found in [33]. $\mathcal{PS}^{-1}$ is the inverse of periodic shuffling operator. $\pi$ and $\pi^{-1}$ are permutation and inverse permutation operations. $Conv_{1}$ and $Conv_{2}$ are convolution kernels, $k_{1}\in\mathbb{R}^{4C\times C/r\times 1\times 1}$, $k_{2}\in\mathbb{R}^{C/r\times 4C\times 1\times 1}$, separately.
To enable the Swin Transformer block to capture both long-range and short-range dependencies, we replace the MLP component with ESGA module, incorporating GELU activation [34]. Fig. 8 shows the detail of Swin Transformer block.
2.2.2 EPGA module
We adapt the ESGA module in a parallel manner at the skip connections in the decoder to enhance channel and spatial interactions while suppressing noise from the encoder, as shown in Fig. 9. Assume the input feature is $F_{cat}\in\mathbb{R}^{N\times H\times W\times 4C}$, the output feature is calculated as below:
$F_{out}=sigmoid(0.5*L_{2}(L_{1}(F_{cat}))+\\ 0.5*\pi^{-1}(\mathcal{PS}(Conv_{2}(Conv_{1}(\pi(\mathcal{PS}^{-1}(F_{cat})))))))\circledast L(F_{cat})$ (3)
where $L$ is MLP layer with weight $W\in\mathbb{R}^{2C\times C}$
2.3 HIC patch expanding module
In Swin-Unet [35], the patch expanding module uses a MLP to expand the channel dimension from $C$ to $4C$, followed by a rearrangement operation, similar to periodic shuffling, to increase the spatial resolution from $H\times W$ to $2H\times 2W$. However, as noted in [27], such learned upsampling strategies often tend to overfit to noise. To address this issue, we propose HIC patch expanding module to force network to learn from low-frequency upsampling interpolation strategies. As illustrated in Fig. 2, our HIC module first uses a MLP for channel expansion. The resulting feature map is then split into two branches: one undergoes nearest neighbor interpolation, while the other applies zero-padding interpolation. A shared convolutional layer is subsequently applied to both branches to ensure the network learns interpolation-invariant features, enhancing robustness and generalization.
Figure 2: HIC patch expanding module. LN represents layer normalization. We adopt nearest interpolation. White boxes represent zero. We expand feature size by interleaving zeros between columns and rows.
2.4 Loss function
We have introduced two kinds of loss functions: Mean Absolute Error (MAE) and the Sobel loss. We choose MAE as loss function, as it can capture high-frequency details [36]. Byeongjoon et al. have compared MAE and MSE loss for CT image denoising[37]. It shows that both MAE and MSE are commonly used, however, both of them fail to capture semantic details. To enhance edge and structural information, we introduce the Sobel loss from [38] that operates in horizontal, vertical, and diagonal directions. Thus, the final loss function is formulated as:
$L_{denoise}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{HW}|X_{LD}^{i}-X_{FD}^{i}|_{1}+\lambda\frac{1}{N}\sum_{i=1}^{N}\frac{1}{HW}|\phi(X_{LD}^{i})-\phi(X_{FD}^{i})|_{2}^{2}$ (4)
where $\phi$ represents the Sobel operator, and the detailed information can be found at [38]. $\lambda$ is the hyperparameter to balance two loss functions.
2.5 Experiments
2.5.1 Dataset and evaluation
The AAPM Low-Dose CT Grand Challenge released a widely used public dataset containing paired LDCT and FDCT images, covering the head, chest, and abdomen. For this study, we selected the abdominal data, where each CT volume has a voxel size of $0.5859\times 0.5859\times 3.0mm^{3}$ and an axial resolution of $512\times 512$. A total of 27 patients were used for training and 9 patients for testing. During preprocessing, the provided DICOM images were first converted to Hounsfield Units (HU), then normalized to the range [0, 1] using a fixed HU window of [-1024, 3072]. The Ultra-Low-Dose PET Imaging (UDPET) Challenge provides a rich and comprehensive dataset designed to advance research in PET image reconstruction under reduced dose settings. PET data used in our work were obtained from the University of Bern, Dept. of Nuclear Medicine and School of Medicine [39]. The dataset comprise 1,447 whole-body 18F-FDG PET scans, acquired across total-body PET systems, Siemens Biograph Vision Quadra: 387 subjects, United Imaging uEXPLORER: 1,060 subjects. From each PET scan, corresponding low-dose images are generated at multiple Dose Reduction Factors (DRFs), specifically at DRFs of 4, 10, 20, 50, and 100, in addition to the original full-dose reference image. Each PET image has an isotropic voxel spacing of 1.65 mm in all three dimensions, with an axial matrix size of 440 $\times$ 440. We randomly pick DRF=100 of 28 patients for training, 9 for testing. We convert the provided DICOM data to standardized uptake values (SUVs) and normalize to the range [0, 1] using min–max scaling. The resulting images were used as input to the network. For quantitative evaluation, evaluated model performance using three standard metrics following the implementation of Chen et al. [10]: peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and root mean squared error (RMSE). For the LDCT dataset, we use the standard deviation as a measure of variability to assess the stability of the models.
2.5.2 Implementation details
We have implemented our model with pytorch 2.7 library and ran the models on two NVIDIA TITAN Xp 12G GPUs. We use Stochastic Gradient Descent (SGD) optimizer with weight decay 1e41e-4. The learning rate decays as follows.
$lr=lr_{base}\cdot(1-\frac{n_{iter}}{M_{total}})^{0.9}$ (5)
where $lr_{base}=0.01$, $n_{iter}$ is the current iterations, $M_{total}$ is the total iterations. We have chosen 3000 epochs to train the model in patches. In training, we set mini-batch to 16, patch size to $64\times 64$. During validation and test, we set batch size to 1, and patch size to $512\times 512$ and $440\times 440$, respectively.
3 Results
3.1 Ablation study
We conduct extensive experiments to evaluate the effectiveness of the proposed method. Specifically, we perform ablation studies on the attention modules and the HIC patch expanding module, with the results presented in Table 1. Our results show that ESGA module is the most important module as the performance drops a lot without it. In addition, we tune the hyperparameter $\lambda$ for the Sobel loss to determine the optimal setting, as shown in Table 2. The study shows that the performance achieves best at $\lambda$ equals to 0.1.
Table 1: Ablation study for HSANet in LDCT dataset. ✗ and ✓ represent whether HSANet include those modules.
ESGA
EPGA
HIC
PSNR 31.9 32.16 32.22 32.13 32.29
Table 2: Effects of factor $\lambda$ for the Sobel loss
$\lambda$ 0 0.1 0.5 1
PSNR 32.28 32.29 32.17 32.01
3.2 Model comparison
We have compared our model with five different methods on LDCT dataset. The compared methods include RED-CNN [10], CTformer [11], UNet [40], Swin-Unet [35] and SwinIR [23].
3.2.1 Quantitative evaluation results on LDCT dataset
We have adopted the same data preparation procedure as the RED-CNN method to ensure a fair comparison. Specifically, pixel intensities were windowed to [-160, 240] and normalized to [0, 1] before evaluation. The average results from 8 independent runs across the 9 test patients are summarized in Table 3. Our method achieves the best performance in terms of all PSNR, SSIM and RMSE among all compared approaches. Moreover, the variance across runs is minimal, much lower than that of RED-CNN, CTformer, and UNet, and comparable to Swin-Unet for PSNR. In terms of RMSE, our method exhibits even less variability than Swin-Unet. Fig. LABEL:LDCT_violin presents violin plots corresponding to the results in Table 3. The plots reveal that the purely CNN-based approach exhibits the largest variance across runs. In contrast, the transformer-only method (CTformer) shows reduced variance, though still greater than the hybrid CNN–transformer approaches. Our method demonstrates the second-highest stability in terms of PSNR and RMSE, being slightly less stable than SwinIR.
Although SwinIR is relatively stable and has a compact model size of 0.41M parameters, its denoising performance is insignificant (PSNR of 31.72), and it requires substantially more GPU memory—approximately 23 GB, which is four times higher than our method. In contrast, our model not only delivers superior denoising accuracy but also demonstrates high stability. Additionally, it is considerably smaller than most competing models and consumes a similar amount of GPU memory, making it well-suited for practical deployment. Fig. LABEL:LDCT_fig presents a visual comparison of pelvis slices from a single patient in the LDCT test set. From the figure, we observe that RED-CNN, Swin-Unet, SwinIR, and CTformer effectively remove background noise, but at the cost of producing smoother images. In contrast, Unet reduces noise while preserving textural details, although some regions around the bone appear grid-like. HSANet overcomes both limitations, suppressing noise while maintaining image sharpness and structural integrity, producing images that closely resemble FDCT.
Table 3: Quantitative evaluation for LDCT dataset. P represents for network parameters
Method PSNR SSIM RMSE GPU P
memory
LDCT 28.84 0.86 15.16 - -
[…]
3.2.2 Quantitative evaluation results on LDPET dataset
We have used the max-min normalization method to normalize images for training. An average test results on all 9 patients are shown in Table 4. As shown in the table, our method outperforms the other approaches and achieves a PSNR of 32.73. Although its performance is 0.31 dB lower than Unet, it achieves a comparable SSIM with substantially fewer parameters. Figure LABEL:LDPET_violin shows the violin plots for the 9 patients in the test set. Our model demonstrates more stable PSNR and RMSE performance across patients compared with Unet. Note that in the PET denoising experiment, we did not use the Sobel loss due to convergence issues.
Figure LABEL:LDPET_fig shows the denoised results of an abdominal image. As observed, CTformer and Swin-Unet produce grid-like artifacts in the image. RED-CNN and SwinIR fail to capture the image patterns. U-Net and HSANet appear to capture them better than the other methods, although their results are overly smooth.
Table 4: Quantitative evaluation for LDPET dataset.
Methods PSNR SSIM RMSE GPU memory parameters
LDPET 27.23 0.91 13.55 - -
RED-CNN 31.79 0.951 7.52 5.08G 1.85M
Swin-Unet 31.36 0.95 7.92 1.2G 0.95M
SwinIR 30.39 0.94 8.9 23G 0.41M
CTformer 31.14 0.92 7.88 4.7G 1.45M
UNet 33.04 0.96 6.72 6G 31M
ours 32.73 0.96 6.9 5.4G 0.61M
4 Discussion and conclusion
To address the challenges posed by complex, structured noise that varies across patients, scanners, and anatomical regions, and to maintain a delicate balance between over-smoothing and under-denoising, which can leave residual artifacts, we developed HASNet, which integrates our proposed efficient attention mechanisms with a novel upsampling strategy, the HIC module. This design leads to state-of-the-art denoising performance, achieving the highest PSNR and lowest RMSE among all compared methods. Beyond accuracy, HASNet demonstrates remarkable stability across multiple runs, with minimal performance variation. Additionally, it is highly efficient: HASNet contains significantly fewer parameters than conventional models and requires substantially less GPU memory, making it not only accurate and robust but also lightweight and practical for real-world clinical deployment.
To summarize, we have developed an efficient denoising model, HSANet. The network is based on Swin-Unet and incorporates an HIC upsampling module to prevent learning from noise, ESA modules to capture semantic structures, and convolutional residual structures to model local dependencies. Our results demonstrate that HSANet is well-suited for clinical use, being stable, lightweight and memory-efficient. In the future, it is worth to explore noise-aware attention architectures that explicitly model the Poisson nature of PET data, enabling more reliable preservation of low-uptake regions.
Acknowledgments
We are grateful for access to the University of Heidelberg’s IWR HPC service for running simulations used in this study.
Appendix A Example Appendix Section
Figure 8: Swin Transformer block enhanced with ESGA
Figure 9: EPGA module. Channel attention and spatial attention with dashed box means without sigmoid activation function